Gain (V/V) R Setting Totals Averages Sample 1 Sample 2 Sample 3 4 ап 7.8 8.1 7.9 3 5.2 6.0 4.3 = 359.3 i=1 j=1 2 4.4 6.9 3.8Use Fishers Least Significant Difference to determine which, if any, pairs of resistance settings significantly affect gainFormulae: an SST = i=1 j=1 -ΣΣνή with an - 1 degrees of freedom SStr = Ś ył_y? with a - 1 degrees of freedom n N i=1 SSE = SSf0.25.01.12 3 4 5 6 7 1 2 3 3 4 1 5.83 2.57 2.02 1.81 1.69 1.62 152 1.57 154 1.54 151 1.SI 140 1.49 1.47 1.46 1.45 1.44 1.430 TABLE. V Percentage Points ta of the t Distribution O. .40 .25 .05 2 .325 .289 .277 .271 .267 1.000 .816 .765 .10 3.078 1.8
Answer #1
Anova: Single Factor
SUMMARY
Groups Count Sum Average Variance
4 3 23.8 7.933333 0.023333
3 3 15.5 5.166667 0.723333
2 3 15.1 5.033333 2.703333
1 3 4.5 1.5 0.39
ANOVA
Source of Variation SS df MS F
Between Groups 62.54917 3 20.84972 21.71846
Within Groups 7.68 8 0.96
Total 70.22917 11

F_{crit}=F_{005;3,8}=4.07

F> Fcrit, so we reject null hypothesis of equal treatments.

LSD=t_{\frac{0.05}{2};9}\sqrt{\frac{2MSE}{3}}=2.262*\sqrt{\frac{2*0.96}{3}}=1.8096

|\bar{x}_{1}-\bar{x}_{2}|= 7.933-5.167 = 2.766 > LSD -> Significant

|\bar{x}_{1}-\bar{x}_{3}|=7.933-5.033=2.9>LSD\Rightarrow Significant

|\bar{x}_{2}-\bar{x}_{3}|=0.134<LSD\Rightarrow not \:significant